Thanks for your help!!!!!!
Solve the systems by graphing.
McGraw Hill 6th edtion. 8.1 # 12
x-2y=8
3x-2y=12

Answer by kev82(151)

Hi,
Problems like this are called simultaneous equations. There are quite a few techniques to solving them. The one I prefer is eliminating one variable by adding.
What we want to do is change one (or both) of the equations so that we can add them together and cancel one varibale.
In this case we have
x - 2y = 8
and
3x - 2y = 12
Let's multiply the first equation by -1. (remember that equations still hold as long as we do the same thing to both sides). We get
-x + 2y = -8
If we now add this with the second equation, we get
(3x - 2y) + (-x + 2y) = 12 + -8
Notice how we have +2y and -2y so all the terms with y have cancelled. We actually have
2x = 4
So x=2
Putting this back into either equation lets us calculate y. Try for yourself, you should get y=-3
Hope that helps,
Kev